OpenStudy (anonymous):

Find the integral of (cos^2(x))((cosxsin^2n+1)*(pi/2*sinx)+(sinxsin^2n+1)(pi/2*cosx)) from 0 to pi/2

OpenStudy (primeralph):

Input it more clearly.

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/2}\cos ^{2}x(\cos x \sin ^{2n+1}(\frac{ \pi }{ 2 }sinx)+sinxsin ^{2n+1}(\frac{ \pi }{ 2 }cosx))dx\]

OpenStudy (primeralph):

Go Wolfram.

OpenStudy (anonymous):

I don't have the first clue on how to use that

OpenStudy (primeralph):

http://integrals.wolfram.com/index.jsp

OpenStudy (swissgirl):

Wolfram sucks these days :(

OpenStudy (primeralph):

How?

OpenStudy (swissgirl):

Thats what happens with all these free awesome programs

OpenStudy (primeralph):

Well, I don't think it even has CAS. I don't use it though.

OpenStudy (swissgirl):

Well if you wanna see the steps then you gotta have membership

OpenStudy (primeralph):

Well, I'd expect them to do that.

OpenStudy (swissgirl):

Ummm this is an annoying function. @Skip2mylou426 what r u currently studying?

OpenStudy (swissgirl):

Wolfram was the best matrix calculator .........

OpenStudy (primeralph):

was?

OpenStudy (primeralph):

Think it's always been Matlab......

OpenStudy (swissgirl):

Welll ya Dont have matlab. I used to have mathematica but my comp broke down so dont have it no more :(

OpenStudy (primeralph):

@Skip2mylou426 I think the integral can't be evaluated easily because of the form sin(acos(b))

OpenStudy (primeralph):

@swissgirl Yeah, if you're into coding, you'll fancy Matlab.

OpenStudy (swissgirl):

I guess he went to the lou :)

OpenStudy (anonymous):

I was trying to type that monstrosity into Wolfram. Epic fail. It didn't work at all.

OpenStudy (anonymous):

http://www.wolframalpha.com/widgets/view.jsp?id=dc816cd78d306d7bda61f6facf5f17f7 This might help too

OpenStudy (swissgirl):

OMG IK THAT

OpenStudy (swissgirl):

It always happens to me

OpenStudy (primeralph):

There is no free CAS engine...........unless you can hack. Otherwise, get your pen and paper.

OpenStudy (swissgirl):

wait let me try wolframming it for u

OpenStudy (anonymous):

A little rusty on the calculus. I haven't taken calc in 3 or 4 years unfortunately

OpenStudy (swissgirl):

What course are u taking now?

OpenStudy (primeralph):

@swissgirl No point. The integrand has the form of a switch variable 'n' which would leave the answer not numeric, but algebraic........which requires a CAS software.

OpenStudy (anonymous):

It's called Problem Solving. It's a combination of every type of math course that could've possibly been taken.

OpenStudy (swissgirl):

FUNNNNNN :)

OpenStudy (anonymous):

haha not quite. He only gives us 1 point per problem solved. All work is required and 500 points are needed to get an A. Class has one month left (halfway done) and I have 178 points right now. KILL ME.

OpenStudy (primeralph):

I don't see the point in putting that into an integral. The integrand can be simplified though....it just looks ugly.

OpenStudy (primeralph):